# UEFA Champions League entry is incentive incompatible again from the 2024/25 season

### v1

@@ Abstract
A tournament is called incentive incompatible if it allows for a situation where a team could be strictly better off by losing. This paper uncovers that the entry rules of the UEFA Champions League, the most prestigious club football competition in Europe, suffer from this shortcoming again from the 2024/25 season. Consequently, a match with misaligned incentives has been barely avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended. Fairness can be easily guaranteed by reversing the order of two paragraphs in the UEFA Champions League regulation.
Keywords: fairness; incentive compatibility; OR in sports; tournament design; UEFA Champions League
@@ 1 Introduction
Incentive compatibility or strategy-proofness—the two phrases are used as synonyms in the following—of a tournament means that each player should have appropriate incentives to perform under any circumstances. Unsurprisingly, incentive compatibility is among the most important tournament design criteria (Devriesere et al.,, 2024). At first glance, losing could never be a dominant strategy in sports. Indeed, if only one tournament is played, every reasonable ranking method satisfies strategy-proofness. However, the rules in the real world are usually more complex and occasionally create perverse incentives for the participants (Csató,, 2021).
According to the seminal paper of Dagaev and Sonin, (2018), the allocation of vacant slots can easily lead to incentive incompatibility. This issue resulted in the violation of strategy-proofness for the UEFA Champions League entry between the 2015/16 and 2017/18 seasons (Csató,, 2019), and the UEFA Champions League seeding from the 2015/16 season (Csató, 2020b, ).
Here we show that, although the recent fundamental reform of the UEFA Champions League has restored the incentive compatibility of the seeding system, the misaligned order in which the vacancies are filled and the novel European performance spots are assigned leads to the violation of strategy-proofness. Furthermore, misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig. However, incentive compatibility can be guaranteed, for instance, by slightly modifying the UEFA Champions League regulation (UEFA,, 2024): the order of applying Paragraphs 3.04 and 3.07 should be reversed. Therefore, UEFA is called to address the worrying deficiency as soon as possible.
The paper is structured as follows. Section 2 gives a concise summary of literature, focusing on the works that uncovered a situation where a team could have been strictly better off by losing. A motivating example is presented in Section 3. Section 4 discusses the root of the problem and offers potential solutions.
@@ 2 Related literature
Since tanking—the act of deliberately losing a game in order to gain some other advantage—is against the spirit of sports, incentive compatibility is an extensively discussed topic in the field of tournament design (Kendall and Lenten,, 2017; Csató,, 2021). Probably the most serious type of incentive incompatibility is when a team can be strictly better off by losing. According to our knowledge, such a possibility has been identified first by Dagaev and Sonin, (2018) in the UEFA Europa League. Several qualification tournaments for the FIFA World Cups and the UEFA European Championships have turned out to suffer from the same weakness (Csató, 2020a, ), even allowing for a match where the dominant strategy of both opponents is to play a draw (Csató, 2020c, ).
@@
Fortunately, Section 4 will reveal that the third problem has been solved by the new seeding rules effective from the 2024/25 season. But the reform has also created potentially misaligned incentives regarding the UEFA Champions League entry as will be seen in Section 3.
Hence, our research is connected to the studies evaluating the effects of rule changes in the UEFA club competitions. Besides Csató, 2020b , the alternative seeding regime of the UEFA Champions League, introduced in the 2015/16 season, has been the subject of two further papers (Corona et al.,, 2019; Dagaev and Rudyak,, 2019). The modified structure of the champions path of the UEFA Champions League qualifying system has been examined by Csató, 2022b . A recent analysis of the away goals rule (Bahamonde-Birke and Bahamonde-Birke,, 2023) provides more insight into the consequences of abolishing this tie-breaking rule from the 2021/22 season. Last but not least, Gyimesi, (2024) uncovers how the new tournament format of the UEFA Champions League from the 2024/25 season affects competitive balance. The current paper is probably this fundamental innovation in the most prestigious European club football competition.
@@ 3 A real-world example
Currently, the four highest-ranked associations—including Germany—are represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders have a guaranteed place in the league phase even if they do not qualify directly through their domestic championship. The possible vacancy is filled according to UEFA, (2024, Paragraph 3.04):
@@
UEFA has introduced the so-called European performance spot from the 2024/25 season (UEFA,, 2024, Paragraph 3.07):
Finally, the two associations whose affiliated clubs achieved the best collective performance in the previous season of the UEFA men’s club competitions (i.e. best season club association coefficient), in accordance with the season coefficient principles (see Annex D.3), are each entitled to an additional place in the league phase, known as a European performance spot (EPS). The EPS is allocated to the club that finishes the relevant association’s domestic championship in the highest position of all those that do not qualify for the league phase of the UEFA Champions League via the domestic championship (after any vacancies have been filled in accordance with Paragraph 3.04 to Paragraph 3.06). The two associations’ overall quota of clubs qualifying to UEFA club competitions are each increased by one club.
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
  • If Borussia Dortmund is the fourth, then the first five teams qualify for the UEFA Champions League independently of the outcome of the 2023/24 UEFA Champions League final. If Dortmund wins the final, a vacancy is created, filled by the Ukrainian champion Shakhtar Donetsk (UEFA,, 2024, Paragraph 3.04). If Dortmund loses the final, this team creates no vacancy.
@@
Although the above pathological situation has not been realised as Borussia Dortmund has been guaranteed to finish in the fifth place before the last round, UEFA has clearly had good luck in avoiding a strange game in the 2023/24 German Bundesliga. Some football fans have noticed the problem of misaligned incentives for Eintracht Frankfurt, too, see https://m.sports.ru/tribuna/blogs/gazzzzzpacho/3239387.html.
@@ 4 Discussion
It should be acknowledged that the UEFA Champions League rules have improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for automatically assigning the titleholder to the first pot (UEFA,, 2023). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@
UEFA, (2024, Paragraph 3.04) has changed compared to the previous season. However, under both rules, no club could have an incentive to create a vacancy in order to obtain the released slot since either the highest-ranked not already qualified club from the association of the titleholder or a club from a different association benefits from the vacancy (see the cumbersome discussion of UEFA, (2024, Paragraph 3.04b)). Consequently, Paragraph 3.04 is not responsible for incentive incompatibility on its own.
On the other hand, the European performance spots are allocated only after the vacancies caused by the two titleholders are solved (UEFA,, 2024, Paragraph 3.07). Thus, a team playing in the same round-robin tournament (domestic championship) as one of the (potential) titleholders can be interested in not creating vacancies in the UEFA Champions League. This can be achieved by “helping” other teams to be ranked over the (potential) titleholder—a clear example of a tanking opportunity as we have seen above. Such misaligned incentives can arise in the associations entitled to at least one spot in the league phase of the UEFA Champions League if they also earn one of the European performance spots. These are currently the 10 highest-ranked associations: in the 2024/25 season, England, Spain, Germany, Italy, France, the Netherlands, Portugal, Belgium, Scotland, and Austria.
Two potential remedies are offered:
  • Rule change A: Modify the order of applying Paragraphs 3.04 and 3.07 in the UEFA Champions League regulation (UEFA,, 2024). That is, first the two European performance spots are allocated (UEFA,, 2024, Paragraph 3.07), and then the possible vacancies created by the titleholders are considered (UEFA,, 2024, Paragraph 3.04).
  • Rule change B: Modify the way of filling vacancies in UEFA, (2024, Paragraph 3.04). That is, any vacancy created by a titleholder is filled by the highest-ranked not already qualified club from its association.
@@
If UEFA is not much worried about manipulation in a domestic league, and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. Following the previous literature, the current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, 2022a ; even if it would be difficult to implement.
@@
Our paper has revealed that the UEFA Champions League rules, similar to three previous editions (Csató,, 2019), are incentive incompatible again from the 2024/25 season. The order in which the vacancies caused by the titleholders are filled and the two European performance spots are allocated has almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt one of our recommendations in order to restore the strategy-proofness of its flagship tournament.
@@ Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838.

### v2

@@ Abstract
A tournament is called incentive incompatible if it allows for a situation where a team could be strictly better off by losing. This paper uncovers that the entry rules of the UEFA Champions League, the most prestigious club football competition in Europe, suffer from this shortcoming again from the 2024/25 season. Consequently, a match with misaligned incentives has been barely avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the UEFA Champions League regulation.
Keywords: fairness; incentive compatibility; integrity; tournament design; UEFA Champions League
@@ 1 Introduction
Incentive compatibility or strategy-proofness—the two phrases are used as synonyms in the following—of a tournament means that each contestant should have appropriate incentives to perform under any circumstances. Unsurprisingly, incentive compatibility is among the most important tournament design criteria (Devriesere et al.,, 2024), and the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante (Scelles et al.,, 2024). At first glance, losing could never be a dominant strategy in sports. Indeed, if only one tournament is played, every reasonable ranking method satisfies strategy-proofness. However, the rules in the real world are usually more complex and occasionally create perverse incentives for the participants (Csató,, 2021).
According to the seminal paper of Dagaev and Sonin, (2018), the allocation of vacant slots can easily lead to incentive incompatibility. Indeed, this issue resulted in the violation of strategy-proofness for the UEFA Europea League entry until the 2014/15 season (Dagaev and Sonin,, 2018), for the UEFA Champions League entry between the 2015/16 and 2017/18 seasons (Csató,, 2019), and the UEFA Champions League seeding from the 2015/16 season (Csató, 2020b, ). The introduction of the UEFA Nations League has also allowed for games where a team has an incentive to lose (Csató,, 2021; Csató, 2022a, ; Haugen and Krumer,, 2021).
Here we show that, although the recent fundamental reform of the UEFA Champions League has restored the incentive compatibility of the seeding system, the misaligned order in which the vacancies are filled and the novel European performance spots are assigned leads to the violation of strategy-proofness. Furthermore, misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig. However, incentive compatibility can be guaranteed, for instance, by slightly modifying the UEFA Champions League regulation (UEFA,, 2024): the order of applying Articles 3.04 and 3.07 should be reversed. Therefore, UEFA is called to address the worrying deficiency as soon as possible.
The paper is structured as follows. Section 2 gives a concise summary of literature, focusing on the works that uncovered a situation where a team could have been strictly better off by losing. A motivating example is presented in Section 3. Section 4 discusses the root of the problem and offers potential solutions, while Section 5 concludes.
@@ 2 Related literature
Since tanking—the act of deliberately losing a game in order to gain some other advantage—is against the spirit of sports, incentive compatibility is an extensively discussed topic in the field of tournament design (Kendall and Lenten,, 2017; Csató,, 2021). Probably the most serious type of incentive incompatibility is when a team can be strictly better off by losing. According to our knowledge, such a possibility has been identified first by Dagaev and Sonin, (2018) in the UEFA Europa League. Several qualification tournaments for the FIFA World Cups and the UEFA European Championships have been demonstrated to suffer from the same weakness (Csató, 2020a, ), even allowing for a match where the dominant strategy of both opponents is to play a draw (Csató, 2020c, ).
@@
Fortunately, Section 4 will reveal that the third problem has been solved by the new seeding rules effective from the 2024/25 season. On the other hand, the reform has also created potentially misaligned incentives regarding the UEFA Champions League entry as will be seen in Section 3.
Hence, our research is connected to the studies evaluating the effects of rule changes in the UEFA club competitions. Besides Csató, 2020b , the alternative seeding regime of the UEFA Champions League, introduced in the 2015/16 season, has been the subject of two further papers (Corona et al.,, 2019; Dagaev and Rudyak,, 2019). The modified structure of the champions path of the UEFA Champions League qualifying system has been examined by Csató, 2022b . A recent analysis of the away goals rule (Bahamonde-Birke and Bahamonde-Birke,, 2023) provides more insight into the consequences of abolishing this tie-breaking rule from the 2021/22 season. Last but not least, Gyimesi, (2024) uncovers how the new tournament format of the UEFA Champions League from the 2024/25 season affects competitive balance. The current paper is among the first ones that address this fundamental innovation in the most prestigious European club football competition.
@@ 3 A real-world example
Currently, the four highest-ranked associations—including Germany—are represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders have a guaranteed place in the league phase even if they do not qualify directly through their domestic championship. The possible vacancy is filled according to UEFA, (2024, Article 3.04):
@@
UEFA has introduced the so-called European performance spot from the 2024/25 season (UEFA,, 2024, Article 3.07):
Finally, the two associations whose affiliated clubs achieved the best collective performance in the previous season of the UEFA men’s club competitions (i.e. best season club association coefficient), in accordance with the season coefficient principles (see Annex D.3), are each entitled to an additional place in the league phase, known as a European performance spot (EPS). The EPS is allocated to the club that finishes the relevant association’s domestic championship in the highest position of all those that do not qualify for the league phase of the UEFA Champions League via the domestic championship (after any vacancies have been filled in accordance with Articles 3.04 to 3.06). The two associations’ overall quota of clubs qualifying to UEFA club competitions are each increased by one club.
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
  • If Borussia Dortmund is the fourth, then the first five teams qualify for the UEFA Champions League independently of the outcome of the 2023/24 UEFA Champions League final. If Dortmund wins the final, a vacancy is created, filled by the Ukrainian champion Shakhtar Donetsk (UEFA,, 2024, Article 3.04). If Dortmund loses the final, this team creates no vacancy.
@@
Although the above pathological situation has not been realised as Borussia Dortmund has been guaranteed to finish in the fifth place before the last round, UEFA has clearly had good luck in avoiding a strange game in the 2023/24 German Bundesliga. Some football fans have also noticed the problem of misaligned incentives for Eintracht Frankfurt, see https://m.sports.ru/tribuna/blogs/gazzzzzpacho/3239387.html.
@@ 4 Discussion
A recent literature survey on the causes and consequences of fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to gather more knowledge on actionable elements with regards to the prevention of fraud. Our paper has examined the new format of the UEFA Champions League from this perspective. It should be acknowledged that the rules have improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for automatically assigning the titleholder to the first pot (UEFA,, 2024, Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@
UEFA, (2024, Article 3.04) has changed compared to the previous season of 2023/24. However, under both rules, no club could have an incentive to create a vacancy in order to obtain the released slot since either the highest-ranked not already qualified club from the association of the titleholder or a club from a different association benefits from the vacancy (see the cumbersome discussion of UEFA, (2024, Article 3.04b)). Consequently, Article 3.04 is not responsible for incentive incompatibility on its own.
On the other hand, the European performance spots are allocated only after the vacancies caused by the two titleholders are solved (UEFA,, 2024, Article 3.07). Thus, a team playing in the same round-robin tournament (domestic championship) as one of the (potential) titleholders can be interested in not creating vacancies in the UEFA Champions League. This can be achieved by “helping” other teams to be ranked over the (potential) titleholder—a clear example of a tanking opportunity as we have seen above. Such misaligned incentives can arise in the associations entitled to at least one spot in the league phase of the UEFA Champions League if they also earn one of the European performance spots. These are currently the 10 highest-ranked associations: in the 2024/25 season, England, Spain, Germany, Italy, France, the Netherlands, Portugal, Belgium, Scotland, and Austria.
We offer the following remedies:
  • Rule change A: Modify the order of applying Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA,, 2024). That is, first the two European performance spots are allocated (UEFA,, 2024, Article 3.07), and then the possible vacancies created by the titleholders are addressed (UEFA,, 2024, Article 3.04).
  • Rule change B: Modify the way of filling vacancies in UEFA, (2024, Article 3.04). That is, any vacancy created by a titleholder is filled by the highest-ranked not already qualified club from its association.
@@
If UEFA is not much worried about manipulation in a domestic league, and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. Following the previous literature, the current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, 2022a , even if it would be difficult to implement.
@@
5 Conclusions
Our paper has revealed that the UEFA Champions League rules, similar to three previous editions (Csató,, 2019), are incentive incompatible again from the 2024/25 season. The order in which the vacancies caused by the titleholders are filled and the two European performance spots are allocated has almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt one of our recommendations in order to safeguard the integrity of its flagship tournament.
@@ Acknowledgements
One anonymous colleague provided valuable comments and suggestions on an earlier draft.
@@
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838, and the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.